# WebDRO: A Web-based Group-level Clustering and Reweighting Method for Unsupervised Dense Retrieval

### v3

@@ 4 Experimental Methodology
This section introduces the training dataset, evaluation, baseline models, and implementation details.
Training and Evaluation Datasets. In our experiments, we use the anchor document pairs from ClueWeb22 Overwijk et al. (2022), a dataset containing extensive web page information, to train dense retrievers in an unsupervised setting. We use MS MARCO Nguyen et al. (2016) and BEIR Thakur et al. (2021), two widely used English retrieval benchmarks, to evaluate the retrieval effectiveness in both unsupervised and finetuned settings. Following previous work Thakur et al. (2021); Xie et al. (2023b), nDCG@10 is used as our evaluation metric.
Implementation Details. Our implementation is based on OpenMatch Liu et al. (2021); Yu et al. (2023). During training the embedding model for clustering, we consider that the URL of web pages contains valuable information to learn the relations among web pages. Therefore, we incorporate the URL with the title and content as the text representation of web pages. We cluster training data into $500$ groups, yet the actual data may contain slightly less than 500 groups because the groups smaller than $128$ are merged. An exception is DRO w/ URL, which contains $4,732$ groups.
For retrieval models, we start from the T5-base Raffel et al. (2020) ($220$M parameters) checkpoint and contrastively train the dense retrieval model on web data with in-batch negatives and BM25 retrieved negatives for 1 epoch. Then we reinitialize the group weight and continuously train the model using hard (more challenging) negatives Xiong et al. (2021) for 1 more epoch. To enhance the stability of WebDRO, we use gradient accumulation and update the group weights every $500$ training step (gradients of model parameters are not cached).
Baseline Models. We compare WebDRO with a commonly used sparse retriever, BM25 Robertson et al. (2009), and five unsupervised dense retrieval models. Anchor-DR Xie et al. (2023b) regards anchor texts as queries and uses anchor-document pairs without clustering. We describe the other four models in Appendix A.2. We implement two variants of our method by using different clustering strategies. DRO w/ URL directly clusters the web pages that share the same domain name into the same group. DRO w/ Anchor-DR uses Anchor-DR to encode the documents for clustering instead of using a task-specifically trained embedding model. These variants are discussed in Sec. 5.2.
We also conduct experiments to assess WebDRO in a fine-tuned setting. We compare three different base models: T5-base, Anchor-DR, and WebDRO. These models are fine-tuned using the training set from MS MARCO Passage for 1 epoch.
@@ 5 Evaluation Results
In this section, we first show the performance of different retrieval models and analyze different clustering methods in our approach. Then we analyze the validity of the group weights learned by WebDRO and compare WebDRO’s performance in different tasks. Finally, some case studies are shown.
@@ 5.1 Overall Performance
Overall, WebDRO outperforms other dense retrieval models by achieving at least a 1.2% improvement on MS MARCO and a 1.1% improvement on the average of BEIR. By utilizing group weights, WebDRO significantly enhances retrieval performance on most datasets, indicating its effectiveness in filtering out noise in anchor-document pairs and pretraining dense retrievers in a robust and efficient way.
As shown in the right part of Table 2, WebDRO significantly outperforms Anchor-DR and T5-base on most datasets when they are all finetuned with MS MARCO data. Additionally, finetuned WebDRO also outperforms BM25 on 10 sub-tasks (of 19) and has an average performance gain of 2.6%. The performance gain of WebDRO further indicates the advantages of our clustering and reweighting method can be brought to the fine-tuned scenarios.
We conduct our experiment five times on WebDRO using different random seeds and report the average score in Table 2. The disparity between the top-performing and lowest-performing models on MS MARCO is less than $0.3\%$. We compute the cosine similarity between the group weight vectors of five WebDRO models and the lowest similarity score observed is $98.968\%$. This illustrates WebDRO’s ability to learn stable group weights, thereby ensuring robust performance and improving generalizability across diverse datasets.
@@ 5.2 Analyses on Clustering Strategy
In the experiment, we conduct experiments to compare the effectiveness of different document clustering strategies, as shown in Table 2.
WebDRO leverages the extensive document relations within the web graph to train an embedding model for clustering. However, DRO w/ Anchor-DR directly employs a dense retrieval model (Anchor-DR) for clustering. As shown in the evaluation results, WebDRO outperforms DRO w/ Anchor-DR by achieving about a 1% improvement. We hypothesize that this disparity arises from the fact that the clustering task concerns the relations between documents, whereas Anchor-DR is trained with anchor-document pairs, emphasizing the modeling of relations between queries and documents. A link predictor style embedding model focuses more on extracting document features to model the relation between web pages, making it more tailored for clustering documents.
URLs are commonly employed to encapsulate the content of a web page, thus DRO w/ URL exhibits an improvement over vanilla Anchor-DR. However, in comparison to DRO w/ URL, WebDRO shows better retrieval performance, illustrating that titles and contents also play a crucial role in enhancing the URL-based representations of web pages. Another reason for DRO w/ URL’s performance decline compared to WebDRO is its coarse clustering for the documents belong to large domains. For instance, all web pages within Wikipedia will be clustered in one single group (this group contains roughly $1$M anchors, constituting $7.2\%$ of whole training data) regardless of their topics.
@@ 5.3 Retrieval Effectiveness in Different Testing Scenarios
This section focuses on evaluating the role of the group weighting mechanism in training dense retrieval models. In Figure 2, we present the performance points of all 18 datasets from BEIR and show the relationship between Anchor-DR’s performance and the relative performance enhancement achieved by WebDRO over Anchor-DR.
The evaluation results show that the performance gap between Anchor-DR and WebDRO tends to be larger when the baseline nDCG@10 is low. It illustrates that WebDRO performs particularly well on difficult retrieval tasks, like scidocs (11.26%), hotpotQA (19.60%) and cqadupstack (14.39%). Conversely, on datasets where the baseline already exhibits strong performance, such as trec-covid (-4.30%) and quora (-2.58%), the model’s enhancements are marginal or even negative. This aligns with the optimization goal of WebDRO, which is that the groups with larger training losses are assigned larger weights to improve the overall performance across all datasets. This phenomenon suggests that WebDRO’s unique optimization strategy effectively identifies and prioritizes challenging tasks during pretraining, offering a strategic advantage for performance gains, particularly in scenarios where traditional models struggle.
5.4 Effectiveness of the Group Weights Learned by WebDRO
In this section, two experiments are conducted to show the effectiveness of the group weights that are learned by WebDRO.
@@ 5.4 Effectiveness of the Group Weights Learned by WebDRO
We first randomly sample 50 groups to track the changes of group weights during training. As seen in Figure 2(a), the color gradient within each group gradually shifts towards darker, indicating an approximate monotonic change in group weights. Then we rank the groups according to the weights that are assigned by our final WebDRO model and select the top-ranked and bottom-ranked 30 groups respectively to observe the trend of weight changes during training. As shown in Figure 2(b), the weights of top-weighted groups and bottom-weighted groups are consistently increased and decreased, further showing the stability of the group weights learned by WebDRO.
Effectiveness of Retrieval Models Trained with Different Weighted Groups. We conduct experiments to pretrain dense retrieval models using groups assigned with different weights derived from WebDRO. Specifically, we select top and bottom 30 groups based on their weight ranks, and then sample 500k anchor-document pairs from top and bottom groups respectively. Subsequently, we utilize these subsets of training data to train dense retrieval models. For comparison, we also train a model using a randomly sampled set of 500k pairs.
@@
Table 3: Performance of retrieval model pretrained with different groups. Models are evaluated with nDCG@10.
As shown in Table 3, the retrieval model trained with top-weighted groups significantly outperforms other models trained with an equal number of 500k anchor-document pairs. It illustrates that these groups that are assigned with high weights contain valuable training signals for training the dense retrievers. The retrieval performance of the model that is trained with randomly sampled training data significantly decreases when we replace the training data with instances in bottom-weighted groups. This highlights the capacity of WebDRO to effectively filter out noise present in the training data.
(a) Top-Weighted Anchors
Anchor: The World Cup Document: In all, 80 nations have played in at least one World Cup. Of these, eight national teams have won the World Cup, and they have added stars to their badges, with each star representing a World Cup victory, with only one exception…
Anchor: Sepolture Premature Document: Have People Been Buried Alive? And could it still happen? Snopes Staff Published 9 June 1999 Image via Shutterstock 10 Claim People have been buried alive by mistake…
(b) Bottom-Weighted Anchors
Anchor: Blood Document: What is an inferior vena cava filter placement? An inferior vena cava filter is a small device that can stop blood clots from traveling through the vena cava vein to the heart and lungs…
Anchor: Evidence Based Physical Therapy Document: Evidence-based physical therapy (EBPT) has been defined as physiotherapy informed by relevant high quality clinical research. The practice of evidence-based physiotherapy should be informed by the integration of…
Table 4: Examples of anchor-document pairs in top-weighted 30 groups and bottom-weighted 30 groups.
@@ 5.5 Analyses of the Loss Landscapes
This experiment shows the WebDRO’s superiority from the perspective of the loss landscape. Loss landscape is a common tool for assessing a model’s stability and generalization capabilities Li et al. (2018); Hao et al. (2019). Specifically, we apply small perturbations to the model’s parameters from two random directions and visualize the losses after perturbations in a plane. The “flatness” of the loss landscape is defined as the size of the region around the origin where the training loss remains low Hochreiter et al. (1997). A flatter loss landscape usually indicates better robustness to noises and a higher chance of converging to the optimal when randomly initialized.
As shown in Figure 4, it is evident that with increasing perturbation, the training loss in WebDRO exhibits a gradual increase, whereas in Anchor-DR, the rise in training loss is steep in specific directions. Besides, it can be observed that WebDRO encompasses a larger area with low training loss (the area inside the red lines) than Anchor-DR. In summary, WebDRO presents a flatter and more uniform landscape than Anchor-DR, indicating enhanced stability and efficiency in training and heightened generalization capabilities.
@@
5.6 Case Study
Finally, we show some case studies in Table 4.
We sample examples from both the top-weighted and bottom-weighted groups. In the top-weighted groups, the majority of anchor-document pairs exhibit the following characteristics: both the anchor and the document contain rich information, and the anchor text serves as the primary topic or theme of the document. A large proportion of anchor-documents pairs in bottom groups exhibit deficiencies that hinder their ability to effectively represent retrieval queries and documents. For instance, in the first example, although the anchor text “blood” is mentioned in the document, it is unrelated to the document’s theme.
We conducted statistical analysis on both the top 30 and bottom 30 groups. As displayed in Table 5, sentences from groups with high weights tend to be longer on average compared to those from groups with low weights. Since functional units (e.g. Learn More, Contact Us) are considered as separate sentences, we speculate that this preference is because functional units will hinder information extraction of the model. Furthermore, our analysis reveals that the model does not exhibit a clear preference between long and short queries. Additionally, the choice of frequent words indicates differing focuses between the top-weighted and bottom-weighted groups. The model appears to favor academic and journalistic content while marginalizing documents of an introductory or advertising nature.
Attributes & Features Top-Weighted Bottom-Weighted
30 Groups 30 Groups
Avg. Query Length 3.43 3.03
Avg. Sentence Length 9.34 8.20
Frequent Words
covid, news,
email, updated,
article, published,
health, secure,
data, privacy
encyc., new,
united, welcome,
university, help,
name, follow,
schools, contact
Table 5: Comparison between top-weighted and bottom-weighted groups. “Avg. Sentence Length” means the average number of words in a sentence in documents. “Frequent Words” are content words that most frequently appear in the corresponding groups.
@@ 6 Conclusion
This paper introduces WebDRO, an unsupervised dense retrieval model that utilizes web structures to cluster documents and adjusts group weights during contrastive training. We use structures of the web graph to train a link prediction style embedding model for clustering and utilize GroupDRO to reweight the clusters during training. Experiments on MS MARCO and BEIR show that WebDRO achieves significant improvements over baseline models, especially on difficult tasks. Further analysis shows that group weights learned by WebDRO are stable and valid, which means that the model’s preferences are consistent throughout training and that groups with larger weights are more beneficial for the model, while noisy or uninformative groups are down-weighted.
@@
In this work, we cluster different contrastive pairs by their semantic features, which means contrastive pairs with similar content will be clustered in the same group. This approach is based on a hypothesis: semantically resembled contrastive pairs have similar contributions to retrieval model training. However, other factors may also influence the data point’s contribution, like format, style, or even features we can’t explicitly describe. In future works, we hope to design a clustering strategy directly related to the optimization goal of dense retrieval to achieve more delicate clusters.
Research has shown that negative documents play a crucial part in dense retrieval, and introducing stronger (more confusing) negatives would enhance the model’s performance Xiong et al. (2021). In our methods, documents within each cluster are similar to each other, thus would serve as stronger negatives. However, to evaluate the impact of the reweighting strategy alone, we didn’t modify the way of choosing negative documents. In future works, we hope to make better use of our clusters and build stronger negatives to achieve better performance.
In this work, we mainly focus on the task of information retrieval itself. As information retrieval has vast applications in many fields, it would be interesting to see WebDRO’s performance in downstream applications like retrieval-augmented generation Guu et al. (2020) and retrieval-augmented diffusion Blattmann et al. (2022).
@@ A.2 Training Details
Sizes of Datasets. The training data for the retrieval model contains $13.8$M contrastive pairs, and the training data for the embedding model contains $5.1$M contrastive pairs. For evaluation benchmarks, MS MARCO contains $8.8$M documents and $8,960$ queries, and BEIR contains $44.0$M documents and $47,282$ queries in total.
@@
Computational Resources. The training of our model uses 16 NVIDIA RTX A6000 GPUs (48GB of memory each). Each epoch of training (for all $13.8$M contrastive pairs) takes 9 to 12 hours depending on the inter-node communication efficiency.
@@ A.3 More Details of Baseline Models
coCondenser Gao and Callan (2022) bases on the Condenser model Gao and Callan (2021) and is trained with token-level and span-level language modeling objectives. Contriever Izacard et al. (2021) constructs the positive pairs with random cropping and inverse cloze task. SPAR Chen et al. (2022) further trains dense retrievers on Wikipedia and aims at imitating a sparse retrieval teacher. AugTriever Meng et al. focuses on extracting the queries for the documents and using them to train retrieval models.

### v4

@@ 4 Experimental Methodology
This section introduces the training dataset, evaluation, baseline models, and details on implementation.
Training and Evaluation Datasets. We use MS MARCO Nguyen et al. (2016) and BEIR Thakur et al. (2021), two widely used English retrieval benchmarks, to evaluate the retrieval effectiveness in both unsupervised and finetuned settings. We collect anchor document pairs from ClueWeb22 Overwijk et al. (2022), a dataset containing extensive web page information, to train dense retrievers in an unsupervised setting. We use the training set of MS MARCO Passage as data for fine-tuning. Following previous work Thakur et al. (2021); Xie et al. (2023b), nDCG@10 is used as our evaluation metric.
Implementation Details. Our implementation is based on OpenMatch Liu et al. (2021); Yu et al. (2023). During training the embedding model for clustering, we think the URL of web pages contains valuable information about the relations between web pages. Therefore, we incorporate the URL with the title and content as the text representation of web pages. We cluster training data into $500$ groups, yet the actual data may contain slightly less than 500 groups because the groups smaller than $128$ are merged (as mentioned in Sec. 3.2). We analyzed the effect of using different group numbers in Appendix A.4.
For unsupervised retrieval models, we start from the T5-base Raffel et al. (2020) (with $220$M parameters) checkpoint and contrastively train the dense retrieval model on ClueWeb data with in-batch negatives and BM25 retrieved negatives for 1 epoch. Then we reinitialize the group weight and train the model for another epoch using in-batch negatives and more challenging negatives generated with ANCE Xiong et al. (2021). To enhance the stability of WebDRO, we use gradient accumulation and update the group weights every $500$ training step (gradients of model parameters are not cached and updated every step).
For the finetuned setting, we tune base models with the training set of MS MARCO Passage with in-batch negatives and BM25 retrieved negatives for 1 epoch. The finetuning process doesn’t involve clustering or reweighting.
Baseline Models. We compare WebDRO with a commonly used sparse retriever, BM25 Robertson et al. (2009), and five unsupervised dense retrieval models. Anchor-DR Xie et al. (2023b) serves as the baseline for direct comparison with WebDRO. Anchor-DR regards anchor texts as queries and uses anchor-document pairs without clustering. For the other four baselines, we describe them in Appendix A.2.
To evaluate WebDRO in a fine-tuned setting, we compare three different base models: T5-base, Anchor-DR, and WebDRO. These base models are finetuned in exactly the same manner.
@@ 5 Evaluation Results
In this section, we first show the performance of different retrieval models and analyze different clustering methods in our approach. Then we compare WebDRO’s performance in different tasks and analyze the validity of the group weights learned by WebDRO.
@@ 5.1 Overall Performance
In the unsupervised setting, WebDRO outperforms Anchor-DR, its counterpart without reweighting by achieving a 1.2% improvement on MS MARCO and a 1.1% improvement on the average of BEIR. By utilizing group weights, WebDRO enhances retrieval performance on most datasets compared with Anchor-DR, indicating its effectiveness in identifying and highlighting rare or challenging data patterns in anchor-document pairs. WebDRO also significantly outperforms the first three unsupervised baselines. Although the amount of data used to train WebDRO is only 5.1% of that used to train E5-Large, WebDRO’s average score only falls behind by 2.1%. This indicates that anchor-document pairs are a high-quality resource for training unsupervised retrievers. We believe stronger dense retrievers can be trained in the same manner as WebDRO after scaling up, and we leave it for future work.
As shown in the right part of Table 1, WebDRO also significantly outperforms Anchor-DR and T5-base on most datasets when they are all finetuned with the same MS MARCO data. Additionally, finetuned WebDRO outperforms BM25 on 10 sub-tasks (out of 19) and has an average performance gain of 2.6%. The performance gain of WebDRO further indicates the advantages of our clustering and reweighting method can be brought to the fine-tuned scenarios.
@@ 5.2 Analyses on Clustering Strategy
To compare the effectiveness of different document clustering strategies, we also implement two variants of WebDRO by using different clustering strategies. DRO w/URL directly clusters the web pages that share the same domain name into the same group. DRO w/Anchor-DR uses Anchor-DR to encode the documents for clustering instead of using a task-specifically trained embedding model. The results in the unsupervised setting are shown in Table 2.
From the table, we can see WebDRO outperforms DRO w/Anchor-DR by achieving about a 1% improvement. We hypothesize that this disparity arises from the fact that the clustering task concerns the relations between documents, whereas Anchor-DR is trained with anchor-document pairs, emphasizing the modeling of relations between queries and documents. A link predictor style embedding model (the one in WebDRO) focuses more on extracting document features to model the relation between web pages, making it more tailored for clustering documents.
URLs are commonly employed to encapsulate the content of a web page, and naturally, DRO w/URL exhibits an improvement over vanilla Anchor-DR. However, in comparison to DRO w/URL, WebDRO shows better retrieval performance, illustrating that titles and contents also play a crucial role in enhancing the URL-based representations of web pages. Another reason for DRO w/URL’s performance decline compared to WebDRO is its coarse clustering for the documents belong to large domains. For instance, all web pages within Wikipedia will be clustered in one single group (this group contains roughly $1$M anchors, constituting $7.2\%$ of whole training data) regardless of their topics.
@@ 5.3 Retrieval Effectiveness in Different Testing Scenarios
This section focuses on evaluating the role of the group weighting mechanism in training dense retrieval models. In Figure 2, we present the performance points of all 18 datasets from BEIR and show the relationship between Anchor-DR’s performance and the relative performance gain achieved by WebDRO over Anchor-DR.
The evaluation results show that the performance gap between Anchor-DR and WebDRO tends to be larger when the baseline nDCG@10 is low. It illustrates that WebDRO performs particularly well on difficult retrieval tasks, like scidocs (with a relative gain of 11.26%), hotpotQA (19.60%) and cqadupstack (14.39%). Conversely, on datasets where the baseline already exhibits strong performance, such as trec-covid (-4.30%) and quora (-2.58%), the model’s enhancements are marginal or even negative. This aligns with the optimization goal of WebDRO, which is that the groups with larger training losses are assigned larger weights to improve the overall performance across all datasets. This phenomenon suggests that WebDRO’s unique optimization strategy effectively identifies and prioritizes challenging tasks during pretraining, offering a strategic advantage for performance gains, particularly in scenarios where traditional models struggle.
5.4 Analyze on the Group Weights Learned by WebDRO
In this section, two experiments are conducted to show the stability and effectiveness of the group weights learned by WebDRO.
@@ 5.4 Analyze on the Group Weights Learned by WebDRO
We first randomly sample 50 groups to track the changes in group weights during training. As seen in Figure 3(a), the color gradient within each group gradually shifts towards darker, indicating an approximate monotonic change in group weights. Then we rank the groups according to the weights that are assigned by our final WebDRO model and select the top-ranked and bottom-ranked 30 groups respectively to observe the trend of weight changes during training. As shown in Figure 3(b), the weights of top-weighted groups and bottom-weighted groups are consistently increased and decreased, further demonstrating the stability of the group weights learned by WebDRO.
Performance of Retrieval Models Trained with Different Weighted Groups. We conduct experiments to pretrain dense retrieval models using groups assigned with different weights derived from WebDRO. Specifically, we select top and bottom 30 groups based on their weight ranks, and then sample 500k anchor-document pairs from top and bottom groups respectively. For comparison, we also train a model using a randomly sampled set of 500k pairs. Subsequently, we utilize these subsets of training data to train dense retrieval models for 1 epoch.
@@
Table 3: Performance of unsupervised retriever trained with different groups. Models are evaluated with nDCG@10.
As shown in Table 3, the unsupervised retriever trained with top-weighted groups significantly outperforms other models trained with an equal number of 500k anchor-document pairs. This illustrates that the groups assigned higher weights are more informative for the model, indicating that they are underrepresented in the training data and post greater challenges. The retrieval performance of the model that is trained with randomly sampled training data significantly decreases when we replace the training data with instances in bottom-weighted groups. This highlights the capacity of WebDRO to effectively identify data already well-interpreted by the model.
@@ 5.5 Analyses of the Loss Landscapes
Loss landscape is a common tool for assessing a model’s stability and generalization capabilities Li et al. (2018); Hao et al. (2019). Specifically, we apply small perturbations to the model’s parameters from two random directions and visualize the losses after perturbations in a plane. The “flatness” of the loss landscape is defined as the size of the region around the origin where the training loss remains low Hochreiter et al. (1997). A flatter loss landscape usually indicates better robustness to noises and a higher chance of converging to the optimal when randomly initialized.
As shown in Figure 4, we can see that with increasing perturbation, the training loss in WebDRO exhibits a gradual increase in all directions, whereas in Anchor-DR, the rise in training loss is steep in specific directions. Besides, it can be observed that WebDRO encompasses a larger area with low training loss (the area inside the red lines) than Anchor-DR. In summary, WebDRO presents a flatter and more uniform landscape than Anchor-DR, indicating enhanced stability in training and heightened generalization capabilities to different tasks.
@@ 6 Conclusion
This paper introduces WebDRO, an unsupervised dense retrieval model that utilizes web structures to cluster documents and adjusts group weights during contrastive training. The training process is divided into two steps. In Step-A, we use structures of the web graph to train a link prediction style embedding model for clustering. In Step-B, we utilize GroupDRO to reweight the clusters during training based on training loss. Experiments on MS MARCO and BEIR show that WebDRO achieves significant improvements over baseline models, especially on difficult tasks. Further analysis shows that group weights learned by WebDRO are stable and valid, which means underrepresented and challenging groups are assigned higher weights consistently throughout training. This confirms WebDRO’s effectiveness in improving dense retrievers’ robustness.
@@
An important part of WebDRO is training the embedding model and generating clusters. However, we only use the cluster information in reweighting, which is a waste of resources. Research has shown that introducing more challenging negatives would enhance the model’s performance Xiong et al. (2021). As being in the same group indicates similarity, we hope to build stronger negatives with clusters to achieve better performance.
In this work, we cluster different contrastive pairs solely by their semantic features, which means contrastive pairs with similar content will be clustered in the same group. This approach is based on a hypothesis: semantically resembled contrastive pairs have similar contributions to retrieval model training. However, other factors may also influence the data point’s contribution, like format, style, or even features we can’t explicitly describe.
Data imbalance in unsupervised training data is the key motivation of our method. Although prior work has confirmed this phenomenon and its consequences, we didn’t closely examine data distribution in our scenario due to our data’s size.
@@ A.2 Training Details
Sizes of Datasets. The training data for training unsupervised the retrieval model contains $13.8$M contrastive pairs, and the training data for the embedding model contains $5.1$M contrastive pairs. For evaluation benchmarks, MS MARCO contains $8.8$M documents and $8,960$ queries, and BEIR contains $44.0$M documents and $47,282$ queries in total.
@@
Computational Resources. The training of our model uses 8 NVIDIA RTX A6000 GPUs (48GB of memory each). Each epoch of dense retrieval model training (for all $13.8$M contrastive pairs) takes 9 to 12 hours depending on the inter-node communication efficiency.
@@ A.3 More Details of Baseline Models
coCondenser Gao and Callan (2022) bases on the Condenser model Gao and Callan (2021) and is trained with token-level and span-level language modeling objectives. SPAR Chen et al. (2022) further trains dense retrievers on Wikipedia and aims at imitating a sparse retrieval teacher. AugTriever Meng et al. focuses on extracting the queries for the documents and using them to train retrieval models. E5-Large Wang et al. (2022) trains retrieval models on CC-Pair, a huge unsupervised dataset collected from multiple sources and curated with a consistency-based filter.